Tamarisk Construction Company uses the percentage-of-completion method of accounting. In 2017, Tamarisk began work under contract #E2-D2, which provided for a contract price of $2,227,000. Other details follow:

2017 2018
Cost incurred during the year $660240 $1422000
Estimated costs to complete, as of December 31 911760 0
Billings to date 425000 2227000
Collections during the year 347000 1523000

What portion of the total contract price would be recognized as revenue in 2017? In 2018?

Revenue recognized in 2017 $______________

Revenue recognized in 2018 $______________

Tamarisk Construction Company uses the percentage-of-completion method of accounting. In 2017, Tamarisk began work under contract #E2-D2, which provided for a contract price of $2,227,000. Other details follow:
Revenue recognized in 2018 $__________?

Prepare a complete set of journal entries for 2017 (using the percentage-of-completion method). (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. For costs incurred use account Materials, Cash, Payables.)

Account Titles Debit Credit
(to record cost of construction)
(to record progress billings)
(to record collections)
(to recognize revenue)

List of accounts

Accounts Receivable
Accounts Payable
Advertising Expense
Allowance for Sales Returns and Allowances
Billings on Construction in Process
Cash
Cash, Parts, Labor
Commission Expense
Commission Revenue
Construction in Process
Construction Expenses
Contract Asset
Contract Liability
Cost of Goods Sold
Cost of Installment Sales
Deferred Gross Profit
Delivery Expense
Discount on Notes Receivable
Estimated Inventory Returns
Finished Goods Inventory
Franchise Revenue
Freight- Out
Gain on Repossession
Income Summary
Installment Accounts Receivable
Installment Sales Revenue
Interest Expense
Interest Revenue
Inventory
Inventory on Consignment
Liability to Bonus Point Customers
Liability to Enyart Company
Liability to Werner Metal Company
License Revenue
Loss from Long-Term Contracts
Loss on Repossession
Materials, Cash, Payables
No Entry
Notes Receivable
Operating Expenses
Payable to Consignor
Purchases
Realized Gross Profit
Repossessed Merchandise
Retained Earnings
Returned Inventory
Revenue from Consignment Sales
Revenue from Franchise Fees
Revenue from Long-Term Contracts
Sales Discounts
Sales Discounts Forfeited
Sales Returns and Allowances
Sales Revenue
Service Revenue
Unearned Franchise Revenue
Unearned Sales Revenue
Unearned Service Revenue
Unearned Warranty Revenue
Warranty Expense
Warranty Liability
Answer #1

Answer

  • All working forms part of the answer
  • Workings

Working

Year end 2017

Year end 2018

A

Contract Price

$       2,227,000.00

$       2,227,000.00

B

Cost Incurred to Date

$           660,240.00

$       1,422,000.00

C

Estimated cost yet to be incurred to complete the contract

$           911,760.00

$                            -  

D = B+C

Total Cost

$       1,572,000.00

$       1,422,000.00

E = (B/D) x 100

% of Completion

42.00%

100.00%

F = A x E

Revenue to date

$           935,340.00

$       2,227,000.00

G

Revenue of Previous year

$                            -  

$          935,340.00

H = F - G

Net Revenue this year

$           935,340.00

$       1,291,660.00

I (=B)

Cost to date

$           660,240.00

$       1,422,000.00

J

Cost to date of previous year

$                            -  

$          660,240.00

K = I - J

Net Cost for the year

$           660,240.00

$          761,760.00

L = H - K

Gross Profits

$           275,100.00

$          529,900.00

  • Answers

--Revenue recognised in 2017 = $ 935,340
--Revenue recognised in 2018 = $ 1,291,660


--Journal Entries

Accounts title

Debit

Credit

Construction in Progress

$               660,240.00

      Cash

$          660,240.00

(construction cost incurred)

Accounts receivables

$               425,000.00

      Billing on Construction contracts

$          425,000.00

(amount billed during the year)

Cash

$               347,000.00

      Accounts receivables

$          347,000.00

(cash received)

Construction in Progress

$               275,100.00

Construction expenses

$               660,240.00

      Construction Revenue

$          935,340.00

(revenue recognised)